OpenStudy (anonymous):

the volume of a sphere is found using the formula v=4/3pir^3 The area of a circle is found using the formula a=pir^2 What is the ratio of the volume to the area?

OpenStudy (anonymous):



OpenStudy (campbell_st):

well almost the ratio is \[\pi r^2 : \frac{4}{3}\pi r^3\] dividing both sides of the ratio by \[\pi r^2\] you will get \[1: \frac{4}{3} r\] or it can be written as \[3:4r\] hope it helps